OpenStudy (gabylovesyou):

what is the x-intercept and y-intercept of the graph of 2x - 4y = 8

OpenStudy (anonymous):

To find the y-intercept make x = 0. To find the x-intercept make y = 0.

OpenStudy (gabylovesyou):

2*0 - 4*0 = 8

OpenStudy (anonymous):

To find y-intercept make x = 0. So: 2x - 4y = 8 2(0) - 4y = 8 Solve. What is y? To find x-intercept make y = 0. So: 2x - 4y = 8 2x - 4(0) = 8 Solve. What is x?

OpenStudy (anonymous):

Is that easier to understand?

OpenStudy (gabylovesyou):

y = -2

OpenStudy (gabylovesyou):

x = 4

OpenStudy (gabylovesyou):

x-intercept is (4,0) and y-intercept is (0,-2)

OpenStudy (anonymous):

Right! :D

OpenStudy (gabylovesyou):

yaay :)